OpenStudy (thefluffmuncher):

Simplify and select the answer with the appropriate restrictions for the variables.

OpenStudy (thefluffmuncher):

\[\frac{ 12x ^{2}-8x }{ 4x ^{2} }\]

OpenStudy (thefluffmuncher):

its say im supposed to set the denominator to zero how do i do this?

OpenStudy (anonymous):

Umm i'm not sure how to do that @dan815 or @jim_thompson5910 might be able to help though Good luck : )

OpenStudy (thefluffmuncher):

thanks

OpenStudy (dan815):

divide each term

OpenStudy (anonymous):

http://www.purplemath.com/modules/rtnldefs.htm

OpenStudy (anonymous):

how do you set it equal to zero?

OpenStudy (anonymous):

write \[4x^2=0\]

OpenStudy (dan815):

3-4/x so when x=0 there is an undefined point

OpenStudy (dan815):

3-2/x*

OpenStudy (thefluffmuncher):

i have to find the restriction

OpenStudy (dan815):

yes so can u see why x=0 is a restriction

OpenStudy (thefluffmuncher):

yea